OpenStudy (anonymous):

Which shows the dimensions of two rectangular prisms that have volumes of 320 mm3 but different surface areas? A. 16 mm by 5 mm by 4 mm; 5 mm by 4 mm by 16 mm B. 10 mm by 4 mm by 6 mm; 6 mm by 10 mm by 4 mm C. 8 mm by 10 mm by 4 mm; 10 mm by 8 mm by 4 mm D. 8 mm by 10 mm by 4 mm; 20 mm by 8 mm by 2 mm

OpenStudy (anonymous):

@e.mccormick @bagajr @kag plz help

OpenStudy (anonymous):

im not sure about this 1

OpenStudy (anonymous):

To find the volume of each one, just multiply the three numbers together. Do you think you can do that for all 8 of them?

OpenStudy (anonymous):

they all show 320 :'(

OpenStudy (anonymous):

Don't be confused; that's just the first half of the problem. Are you sure all of them are 320? Try calculating it again for the prisms in B.

OpenStudy (anonymous):

lol messed tht 1 up its 240

OpenStudy (anonymous):

Good, so you eliminated B.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Now for the second half. The surface area of a rectangular prism (where the dimensions are x, y, and z) is:\[\Large2xy +2xz + 2yz\]

OpenStudy (anonymous):

:0 letters..i never understand those

OpenStudy (anonymous):

The first number is x, the second is y, and the third is z. What I wrote up there means (2 times x times y) plus (2 times x times z) plus (2 times y times z). For example, the surface area of the first one in A is\[\Large(2 \times x \times y) + (2 \times x \times z) + (2 \times y \times z)\]\[\Large (2 \times 16 \times 5)+(2 \times 16 \times 4)+(2 \times 5 \times 4)\]\[\Large160+128+40\]\[\Large328\]

OpenStudy (anonymous):

656?...whats tht about

OpenStudy (anonymous):

but i still dont get it how do i get the answer to this

OpenStudy (anonymous):

plz just can u explain this a little easier

OpenStudy (anonymous):

Sorry, I had to leave for a couple minutes. Do you get what I did for A?

OpenStudy (anonymous):

^look above..

OpenStudy (anonymous):

Do you get how I got the surface area for the first one?

OpenStudy (anonymous):

no :'( im just stupid then im never gonna get this right

OpenStudy (anonymous):

Do you know how to get the area of each face of the prism?

OpenStudy (anonymous):

no thts why im asking for help ive been on this question for 45 min

OpenStudy (anonymous):

30*

OpenStudy (wolf1728):

Look at the original problem. Notice that in 3 of the 4 answers the numbers are the same except in choice D.

OpenStudy (anonymous):

wow...ty for tht

OpenStudy (wolf1728):

Thanks :-)

OpenStudy (anonymous):

so are u saying its d?

OpenStudy (anonymous):

Sorry it keeps logging me out and disconnecting me for some reason. Thanks to wolf1728's observation, I guess calculations aren't actually necessary, lol. Yes, it would be D.

OpenStudy (anonymous):

ty both of u

OpenStudy (anonymous):

ok imma bout to post my answers i have to get an 80 to pass ill tell u what i get in a second

OpenStudy (anonymous):

i got A 100 WOOOOOOHOOOOOO

OpenStudy (wolf1728):

By the way to save you some time, here's a calculator that gives volume and surface area at the same time - http://1728.org/recsolid.htm

OpenStudy (anonymous):

wow ty ill bookmark it

OpenStudy (smartyprincess):

me too! ty!